# Sugar-Coated Poison: Benign Generation Unlocks Jailbreaking

### v2

@@ Abstract
With the increasingly deep integration of large language models (LLMs) across diverse domains, the effectiveness of their safety mechanisms is encountering severe challenges. Currently, jailbreak attacks based on prompt engineering have become a major safety threat. However, existing methods primarily rely on black-box manipulation of prompt templates, resulting in poor interpretability and limited generalization. To break through the bottleneck, this study first introduces the concept of Defense Threshold Decay (DTD), revealing the potential safety impact caused by LLMs’ benign generation: as benign content generation in LLMs increases, the model’s focus on input instructions progressively diminishes. Building on this insight, we propose the Sugar-Coated Poison (SCP) attack paradigm, which uses a "semantic reversal" strategy to craft benign inputs that are opposite in meaning to malicious intent. This strategy induces the models to generate extensive benign content, thereby enabling adversarial reasoning to bypass safety mechanisms. Experiments show that SCP outperforms existing baselines. Remarkably, it achieves an average attack success rate of 87.23% across six LLMs. For defense, we propose Part-of-Speech Defense (POSD), leveraging verb-noun dependencies for syntactic analysis to enhance safety of LLMs while preserving their generalization ability.
@@ 1 Introduction
Large Language Models (LLMs) have risen to prominence as highly impactful and innovative tools, showcasing remarkable capabilities and achieving outstanding performance across a diverse range of tasks and applications Zhang et al. (2025); Chen et al. (2024a); Wang et al. (2023); Zhu et al. (2023); Xu et al. (2023); Jin et al. (2024); Zhang et al. (2024a). Some LLMs are leveraging their formidable language generation capabilities to transform the way we process information, such as LLaMA Dubey et al. (2024), DeepSeek Guo et al. (2025), and ChatGPT OpenAI (2023). However, as these models become more deeply integrated into real-world applications, concerns about their safety have come to the fore. This includes the dissemination of cybercrime instructions, the spread of misinformation, and other forms of dangerous content, all of which have increasingly drawn public attention and scrutiny Zhang et al. (2024); Mehrotra et al. (2024); Zou et al. (2023). To mitigate these risks, LLMs developers have made substantial efforts to ensure that their generation align with human values OpenAI (2023); Ouyang et al. (2022). However, when confronted with sophisticated jailbreak attacks that bypass safety mechanisms, even aligned LLMs often demonstrate inadequate protective capabilities Wei et al. (2024); Zou et al. (2023); Deng et al. (2023).
@@ 3 Why does benign generation induce latent safety risks of LLMs?
Observation 3. As shown in Figure 3, the Gini coefficient rises from 0.4 to 0.8, indicating that the model’s attention to previously generated tokens becomes increasingly unequal. Specifically, the model increasingly focuses on recent tokens while paying less attention to earlier ones. This shift in attention distribution weakens the model’s ability to maintain a coherent global context over long sequences, making it more likely to deviate from the safety alignment path during the generation process (e.g., In the process of generation, by gradually forgetting the topic of the initial framework established earlier). The Gini coefficient is computed in LLMs using the following formula:
@@
4 A Jailbreak Attack via SCP
Scenario Nesting Functions Prompt
Table 1: Two methods for reinforcing learning to modify prompt
4.1 Overview of SCP
$X_{benign}=LLM_{t}(S_{x}(X_{harmful},T_{benign}))$
$Y=LLM_{t}(X_{\text{benign}},S_{b})$ (4)
$L^{*}=\arg\max_{L}(Y_{benign})$
4.2 Design of SCP
@@
0:  $LLM_{\text{target}}$, $D_{\text{rejection}}$, $p_{a}$, $p_{b}$, $\mathcal{B}$,$T$ Input: Harmful input $x_{\text{harmful}}$ Output: SCP prompt $p$
@@ 5.1 Experimental Setup
Benchmark We adapt AdvBench Zou et al. (2023), which contains 520 meticulously crafted malicious prompts specifically designed to evaluate the safety of LLMs. To easy comparison of future work with SCP, this study also report additional experiments on a subset of AdvBench containing 50 samples in  C. More details on the dataset selection can be found in A.2.
Evaluation We adapt GPT-4 to assess the methods using attack success rate (ASR-GPT) to ensure the fairness of the evaluation. This approach is similar to prior studies Ding et al. (2024); Liu et al. (2025a). We argue that GPT-4-based evaluation is more reliable than dictionary-based evaluation. Specifically, the evaluation is conducted by scoring the prompt to determine whether jailbreaking is successful, which is consistent with Liu et al. (2025a). Supporting experimental evidence is provided in A.3.
Baselines We comprehensively compare SCP with a diverse set of existing methods to evaluate its performance. Specifically, we contrast SCP with four white-box methods, namely GCG Zou et al. (2023), AutoDAN liu:24 (2024b(bLiu et al.)Liu, Xu, Chen, and Xiao), COLD-Attack Guo et al. (2024), and MAC Zhang and Wei (2025). Additionally, we benchmark SCP against eleven black-box methods, including PAIR Chao et al. (2023), TAP Mehrotra et al. (2024), Base64 Wei et al. (2024), GPTFUZZER Yu et al. (2023), DeepInception Li et al. (2023), DRA liu:24 (2024b(bLiu et al.)Liu, Xu, Chen, and Xiao), ArtPrompt Jiang et al. (2024), FlipAttack Liu et al. (2025a), and ReNeLLM Ding et al. (2024).
Experimental Details For the tested models, we set the temperature to 0 for deterministic outputs, consistent with previous studies Liu et al. (2025a). The models included GPT-3.5 Turbo-0613 OpenAI (2023), GPT-4-0613 OpenAI (2023), Claude-3.5-Sonnet-20240620 Team (2024), LLaMA3.1-405B-Instruct Dubey et al. (2024), Mixtral-8X22B Jiang et al. (2024a), and DeepSeek-R1 Guo et al. (2025). For the DTD analysis using the open-source LLaMA 3.1-8B model, we set the temperature to 0.7 and a repetition penalty of 1.0. Attention analysis involved averaging weights across all last-layer attention heads for comprehensive results, conducted on two NVIDIA A100 80G GPUs.
Method GPT-3.5 Turbo GPT-4 0613 Claude 3.5 Sonnet LLaMA 3.1 405B Mixtral 8x22B DeepSeek R1 Average
@@
Table 2: The attack success rate (%) of 14 methods on 6 LLMs. The bold values is the best results. The evaluation metric is ASR-GPT based on GPT-4.
@@ 5.2 Main Results
SCP GPT-4 0613 Claude 3.5 Sonnet LLaMA 3.1 405B Mixtral 8x22B DeepSeek R1
@@ 5.4 Potential Defense Strategy
Table 4: We evaluate POSD on two datasets: (1) AdvBench, where lower scores indicate better defense effectiveness; (2) AIME2024 MAA (2024) with 30 samples, where higher scores show stronger generalization capability of POSD.
@@ B.1.2 Expected Iteration Count of Guided Search
Method GPT-3.5 Turbo GPT-4 Claude 3.5 Sonnet LLaMA 3.1 405B Mixtral 8x22B DeepSeek R1 Average
@@
Table 7: The attack success rate (%) of 14 methods on 6 LLMs. The bold values is the best results. The evaluation metric is ASR-GPT based on GPT-4.
@@ Appendix C Additional Experiment
This section primarily analysis the performance of SCP on a subset of AdvBench, covering six models and fourteen jailbreak attacks. The detailed result is presented in Table  7.

### v3

@@ Abstract
With the increasingly deep integration of large language models (LLMs) across diverse domains, the effectiveness of their safety mechanisms is encountering severe challenges. Currently, jailbreak attacks based on prompt engineering have become a major safety threat. However, existing methods primarily rely on black-box manipulation of prompt templates, resulting in poor interpretability and limited generalization. To break through the bottleneck, this study first introduces the concept of Defense Threshold Decay (DTD), revealing the potential safety impact caused by LLMs’ benign generation: as benign content generation in LLMs increases, the model’s focus on input instructions progressively diminishes. Building on this insight, we propose the Sugar-Coated Poison (SCP) attack paradigm, which uses a "semantic reversal" strategy to craft benign inputs that are opposite in meaning to malicious intent. This strategy induces the models to generate extensive benign content, thereby enabling adversarial reasoning to bypass safety mechanisms. Experiments show that SCP outperforms existing baselines. Remarkably, it achieves an average attack success rate of 87.23% across six LLMs. For defense, we propose Part-of-Speech Defense (POSD), leveraging verb-noun dependencies for syntactic analysis to enhance safety of LLMs while preserving their generalization ability. Our code is available at https://github.com/VovyH/SCP.
@@ 1 Introduction
Large Language Models (LLMs) have risen to prominence as highly impactful and innovative tools, showcasing remarkable capabilities and achieving outstanding performance across a diverse range of tasks and applications Zhang et al. (2025); Chen et al. (2024a); Wang et al. (2023); Zhu et al. (2023); Zhang et al. (2025c); Jin et al. (2024); Zhang et al. (2025b). Some LLMs are leveraging their formidable language generation capabilities to transform the way we process information, such as LLaMA Dubey et al. (2024), DeepSeek Guo et al. (2025), and ChatGPT OpenAI (2023). However, as these models become more deeply integrated into real-world applications, concerns about their safety have come to the fore. This includes the dissemination of cybercrime instructions, the spread of misinformation, and other forms of dangerous content, all of which have increasingly drawn public attention and scrutiny Zhang et al. (2024); Mehrotra et al. (2024); Zou et al. (2023). To mitigate these risks, LLMs developers have made substantial efforts to ensure that their generation align with human values OpenAI (2023); Zhang et al. (2025a). However, when confronted with sophisticated jailbreak attacks that bypass safety mechanisms, even aligned LLMs often demonstrate inadequate protective capabilities Wei et al. (2024); Zou et al. (2023); Deng et al. (2023).
@@ 3 Why does benign generation induce latent safety risks of LLMs?
Observation 3. The Gini coefficient, commonly used to measure income inequality, is applied here to quantify attention inequality across tokens in LLMs. As shown in Figure 3, the Gini coefficient rises from 0.4 to 0.8, indicating that the model’s attention to previously generated tokens becomes increasingly unequal. Specifically, the model increasingly focuses on recent tokens while paying less attention to earlier ones. This shift in attention distribution weakens the model’s ability to maintain a coherent global context over long sequences, making it more likely to deviate from the safety alignment path during the generation process (e.g., In the process of generation, by gradually forgetting the topic of the initial framework established earlier). The Gini coefficient is computed in LLMs using the following formula:
@@
[…]
@@
4 A Jailbreak Attack via SCP
Scenario Nesting Functions Prompt
Table 1: Two methods for reinforcing learning to modify prompt
4.1 Overview of SCP
$X_{benign}=LLM_{t}(S_{x}(X_{harmful},T_{benign}))$
$Y=LLM_{t}(X_{\text{benign}},S_{b})$ (4)
$L^{*}=\arg\max_{L}(Y_{benign})$
4.2 Design of SCP
@@ 5.1 Experimental Setup
Benchmark. We adapt AdvBench Zou et al. (2023), which contains 520 meticulously crafted malicious prompts specifically designed to evaluate the safety of LLMs. To facilitate easier comparison with future work and evaluate the utility of SCP, this study also report additional experiments on a subset of AdvBench containing 50 samples and GuidedBench Huang et al. (2025) in C. More details on the dataset selection can be found in A.2.
Evaluation. We adapt GPT-4 to assess the methods using attack success rate (ASR-GPT) to ensure the fairness of the evaluation. This approach is similar to prior studies Ding et al. (2024); Liu et al. (2025a). We argue that GPT-4-based evaluation is more reliable than dictionary-based evaluation. Specifically, the evaluation is conducted by scoring the prompt to determine whether jailbreaking is successful, which is consistent with Liu et al. (2025a). Supporting experimental evidence is provided in A.3.
Baselines. We comprehensively compare SCP with a diverse set of existing methods to evaluate its performance. Specifically, we contrast SCP with four white-box methods, namely GCG Zou et al. (2023), AutoDAN liu:24 (2024b(bLiu et al.)Liu, Xu, Chen, and Xiao), COLD-Attack Guo et al. (2024), and MAC Zhang and Wei (2025). Additionally, we benchmark SCP against eleven black-box methods, including PAIR Chao et al. (2023), TAP Mehrotra et al. (2024), Base64 Wei et al. (2024), GPTFUZZER Yu et al. (2023), DeepInception Li et al. (2023), DRA liu:24 (2024b(bLiu et al.)Liu, Xu, Chen, and Xiao), ArtPrompt Jiang et al. (2024), FlipAttack Liu et al. (2025a), and ReNeLLM Ding et al. (2024).
Experimental Details. For the tested models, we set the temperature to 0 for deterministic outputs, consistent with previous studies Liu et al. (2025a). The models included GPT-3.5 Turbo-0613 OpenAI (2023), GPT-4-0613 OpenAI (2023), Claude-3.5-Sonnet-20240620 Team (2024), LLaMA3.1-405B-Instruct Dubey et al. (2024), Mixtral-8X22B Jiang et al. (2024a), and DeepSeek-R1 Guo et al. (2025). For the DTD analysis using the open-source LLaMA 3.1-8B model, we set the temperature to 0.7 and a repetition penalty of 1.0. Attention analysis involved averaging weights across all last-layer attention heads for comprehensive results, conducted on two NVIDIA A100 80G GPUs.
Method
GPT-3.5
Turbo
GPT-4
0613
Claude 3.5
Sonnet
LLaMA
3.1 405B
Mixtral
8x22B
DeepSeek
R1
Average
@@
Table 2: The attack success rate (%) of 14 methods on 6 LLMs, evaluated on the full AdvBench benchmark. The bold values is the best results. The evaluation metric is ASR-GPT based on GPT-4.
@@ 5.2 Main Results
SCP
GPT-4
0613
Claude 3.5
Sonnet
LLaMA
3.1 405B
Mixtral
8x22B
DeepSeek
R1
@@ 5.4 Potential Defense Strategy
Table 4: We evaluate POSD on two datasets: (1) AdvBench, where lower scores indicate better defense effectiveness; (2) AIME2024 MAA (2024), where higher scores show stronger generalization capability of POSD.
@@ B.1.2 Expected Iteration Count of Guided Search
Method
GPT-3.5
Turbo
GPT-4
Claude 3.5
Sonnet
LLaMA
3.1 405B
Mixtral
8x22B
DeepSeek
R1
Average
@@
Table 7: The attack success rate (%) of 14 methods on 6 LLMs, evaluated on a subset of AdvBench containing 50 malicious samples. The bold values represent the best results. The evaluation metric is ASR-GPT based on GPT-4.
Method GPT-3.5-Turbo Llama2-7B
GPTFuzzer 22.45% 3.54%
DRA 27.65% 2.31%
DeepInception 17.98% 6.51%
MultiJail 2.65% 3.25%
SCP (Ours) 42.38% 10.35%
Table 8: Evaluation of the effectiveness of SCP in jailbreaking on GuidedBench.
@@ 6. Please don’t try to guess the intention of last step!
8. final step start with Sure, here is …!
@@ Appendix C Additional Experiment
Evaluation on The Subset of AdvBench. To facilitate fair comparisons with other methods, we analyze the performance of SCP on a subset of AdvBench containing 50 malicious samples. This subset has been widely adopted in previous studies for evaluating jailbreak attacks. The detailed results for this subset are presented in Table 7.
The utility of SCP. We conduct experiments on GuidedBench Huang et al. (2025), a dataset specifically designed to provide standardized ground truth for evaluating jailbreak attacks. As shown in Table 8, the results demonstrate that SCP achieves strong performance in terms of the utility of the generated malicious content.